# Inverse design of multiresonance filters via quasi-normal mode theory

### v1

@@ V.2 $LC$-ladder electronic filter
Using the 5th-order circuit of Fig. 6, we optimized the $L$ and $C$ values for our poles criteria Eqs. (5) to obtain Chebyshev bandpass filters, again with 1% bandwidth around $\omega=1$ and minimum $-0.25$dB transmission in the passband. The criteria Eq. (6) to set the background $C(\omega)\rightarrow 0$ were not needed. Our starting element values were chosen as $100$ for each series $L$ and parallel $C$, and $0.01$ for each series $C$ and parallel $L$, with the motivation that initially each series and parallel branch is resonant at $11/\sqrt{L_{k}C_{k}}=\omega=$ 1 with $Q=\omega L^{\text{s}}_{k}/R_{\text{l}}$ or $\omega C^{\text{p}}_{k}/R_{\text{l}}=100=1/$bandwidth. For a 5th-order filter, with the standard $R_{\text{l}}=R_{\text{g}}$ and $\tilde{\sigma}_{1}=-1$, the result was, unsurprisingly, the textbook analytical solution. Then, we optimized for a 4th-order filter, with $R_{\text{l}}/R_{\text{g}}=2A-1+2\sqrt{A(A-1)}=1.6196$ where $A=10^{0.25/10}$ [26] and $\tilde{\sigma}_{1}=-i$, and again we obtained the standard 4th-order solution with $L_{5}=0$ and $C_{5}\rightarrow\infty$ (see Fig. 6 caption): the optimization “discovered” that it could obtain a 4th-order solution by effectively removing part of the circuit. However, our approach also allows us to seek more non-standard solutions. In particular, we targeted the same 4th-order Chebyshev filter, but with the non-standard phase $\tilde{\sigma}_{1}=1$. We find a new solution (values in the Fig. 6 caption indicate $L_{4}$ is removed) that matches the desired transmission amplitude spectrum with an error of at most $0.5\%$ in $|S_{21}|^{2}$, but with an additional constant $\pi/2$ phase shift over the entire frequency range. Such solutions could be very useful, allowing one to combine filtering and phase-shifting functionalities.
[…]
@@ Appendix B Underdetermined Levenberg-Marquardt algorithm
Recall that to minimize $||\boldsymbol{f}(\boldsymbol{x})||^{2}$, the Gauss–Newton method at the $i$-th iteration finds the update $\boldsymbol{x}_{i}+\boldsymbol{\delta}_{i}$ to the current value $\boldsymbol{x}_{i}$ by minimizing the residual under the first-order approximation $||\boldsymbol{f}(\boldsymbol{x}_{i})+\boldsymbol{J}(\boldsymbol{x}_{i})\boldsymbol{\delta}_{i}||^{2}$ (where $\boldsymbol{J}(\boldsymbol{x})=\frac{\partial\boldsymbol{f}}{\partial\boldsymbol{x}}$ is the Jacobian), whose solution for $\boldsymbol{\delta}_{i}$ can be found using the standard normal equation of the ordinary least square:
@@
$[\boldsymbol{J}(\boldsymbol{x}_{i})^{T}\boldsymbol{J}(\boldsymbol{x}_{i})\boldsymbol{+}\lambda_{i}\boldsymbol{\Delta}]\cdot\boldsymbol{\delta}_{i}=-\boldsymbol{J}(\boldsymbol{x}_{i})^{T}\boldsymbol{f}(\boldsymbol{x}_{i})$ (18)
@@
Usually, the nonlinear least-square problems involve more equations than the number of variables (#rows of $\boldsymbol{J}\gg$columns of $\boldsymbol{J}$), so the system is overdetermined. However, in the case of TopOpt, there are usually many more variables than equations (#rows of $\boldsymbol{J}\ll$columns of $\boldsymbol{J}$), making the system underdetermined. Therefore, the matrix $\boldsymbol{J}^{T}\boldsymbol{J}$ is large and rank-deficient, making the regular LMA formulation of Eq. (18) expensive and sometimes (when $\lambda_{i}\ll 1$) even unstable to solve.
@@ C.1 Dielectric 1D-layered stack [72]
$\begin{split}T^{(k)}=&\begin{pmatrix}e^{-ik_{o}n_{k}d_{k}}&0\\ 0&e^{ik_{o}n_{k}d_{k}}\end{pmatrix}\\ T^{(k,k+1)}=&\frac{1}{2\sqrt{\rho_{k,k+1}}}\begin{pmatrix}1+\rho_{k,k+1}&1-\rho_{k,k+1}\\ 1-\rho_{k,k+1}&1+\rho_{k,k+1}\end{pmatrix},\end{split}$ (21)
@@
To calculate the gradient with respect to the layer thicknesses $\{d_{k}\}$, it is easy to calculate $dT/dd_{k}$, as only the element $T^{(k)}$ in Eq. (22) depends on $d_{k}$. Note that, in this case of transfer matrices, the adjoint method for calculating the derivative is essentially equivalent to saving all intermediate matrices $T^{(1...k-1)}$ during the forward $k=1...K$ matrix multiplications in Eq. (22), and then calculating $dT/dd_{k}$ backwards for $k=K...1$ and evaluating only once each matrix $T^{(k+1...K)}$ [81].
@@ C.2 Electronic $LC$-ladder circuit [73]
The $\mathcal{ABCD}$ matrix $\mathcal{T}$ of a two-port electronic circuit is defined as $\begin{pmatrix}v_{1}\\ i_{1}\end{pmatrix}=\begin{pmatrix}\mathcal{A}&\mathcal{B}\\ \mathcal{C}&\mathcal{D}\end{pmatrix}\begin{pmatrix}v_{2}\\ -i_{2}\end{pmatrix}$. Note that, although this $\mathcal{T}$ matrix relates voltage/current at port 1 to those at port 2, thus can also be cascaded, it is not the system transfer matrix $T$, which relates power-normalized inputs/outputs and requires knowledge of the input/output impedances. For an $LC$-ladder circuit, consisting of alternating series $L$$C$ and shunt parallel $L\parallel C$ sections, the respective $\mathcal{ABCD}$ matrices are
$\begin{split}\mathcal{T}^{\text{s}(k)}=&\begin{pmatrix}1&L_{k}s+1/C_{k}s\\ 0&1\end{pmatrix}\\ \mathcal{T}^{\text{p}(k)}=&\begin{pmatrix}1&0\\ C_{k}s+1/L_{k}s&1\end{pmatrix},\end{split}$ (24)

### v2

@@ V.2 $LC$-ladder electronic filter
Using the 5th-order circuit of Fig. 6, we optimized the $L$ and $C$ values for our poles criteria Eqs. (5) to obtain Chebyshev bandpass filters, again with 1% bandwidth around $\omega=1$ and minimum 0.25-0.25dB transmission in the passband. The criteria Eq. (6) to set the background $C(\omega)\rightarrow 0$ were not needed. Our starting element values were chosen as $100$ for each series $L$ and parallel $C$, and $0.01$ for each series $C$ and parallel $L$, with the motivation that initially each series and parallel branch is resonant at $1/\sqrt{L_{k}C_{k}}=\omega=1$ with $Q=\omega L^{\text{s}}_{k}/R_{\text{l}}$ or $\omega C^{\text{p}}_{k}/R_{\text{l}}=100=1/$bandwidth. For a 5th-order filter, with the standard $R_{\text{l}}=R_{\text{g}}$ and $\tilde{\sigma}_{1}=-1$, the result was, unsurprisingly, the textbook analytical solution. Then, we optimized for a 4th-order filter, with $R_{\text{l}}/R_{\text{g}}=2A-1+2\sqrt{A(A-1)}=1.6196$ where $A=10^{0.25/10}$ [20] and $\tilde{\sigma}_{1}=-i$, and again we obtained the standard 4th-order solution with $L_{5}=0$ and $C_{5}\rightarrow\infty$ (see Fig. 6 caption): the optimization “discovered” that it could obtain a 4th-order solution by effectively removing part of the circuit. However, our approach also allows us to seek more non-standard solutions. In particular, we targeted the same 4th-order Chebyshev filter, but with the non-standard phase $\tilde{\sigma}_{1}=1$. We find a new solution (values in the Fig. 6 caption indicate $L_{4}$ is removed) that matches the desired transmission amplitude spectrum with an error of at most $0.5\%$ in $|S_{21}|^{2}$, but with an additional constant $\pi/2$ phase shift over the entire frequency range. Such solutions could be very useful, allowing one to combine filtering and phase-shifting functionalities.
Figure 6: (a) 5th-order LCLC-ladder electrical circuit between a generator with impedance $R_{\text{g}}$ and a load $R_{\text{l}}$, with a series branch first. (b) Transmission phase for two designed 4th-order Chebyshev bandpass filters: the (textbook) standard one with $\tilde{\sigma}_{1}=-i$ [using $L_{k}=(137.8,7.878\text{m},205.6,11.75\text{m},0)$, $C_{k}=(7.256\text{m},$ $126.9,$ $4.864\text{m},$ $85.10,$ $\infty)$; $\text{m}=10^{-3}$] is easily found by our optimization method, but another non-standard optimization solution with $\tilde{\sigma}_{1}=1$ [using $L_{k}=(137.3,$ $7.808\text{m},$ $202.0,$ $\infty,$ $222.8)$, $C_{k}=(7.282\text{m},$ $128.1,$ $4.991\text{m},$ $0.624,$ $4.521\text{m})$] exhibits a constant-$\pi/2$ phase-shifted response.
@@ Appendix B Underdetermined Levenberg-Marquardt algorithm
Recall that to minimize $\boldsymbol{f}(\boldsymbol{x})||^{2}$, the Gauss–Newton method at the $i$-th iteration finds the update $\boldsymbol{x}_{i}+\boldsymbol{\delta}_{i}$ to the current value $\boldsymbol{x}_{i}$ by minimizing the residual under the first-order approximation $\boldsymbol{f}(\boldsymbol{x}_{i})+\boldsymbol{J}(\boldsymbol{x}_{i})\boldsymbol{\delta}_{i}||^{2}$ (where $\boldsymbol{J}(\boldsymbol{x})=\frac{\partial\boldsymbol{f}}{\partial\boldsymbol{x}}$ is the Jacobian), whose solution for $\boldsymbol{\delta}_{i}$ can be found using the standard normal equation of the ordinary least square:
@@
$\boldsymbol{J}(\boldsymbol{x}_{i})^{T}\boldsymbol{J}(\boldsymbol{x}_{i})\boldsymbol{+}\lambda_{i}\boldsymbol{\Delta}]\cdot\boldsymbol{\delta}_{i}=-\boldsymbol{J}(\boldsymbol{x}_{i})^{T}\boldsymbol{f}(\boldsymbol{x}_{i})$ (18)
@@
Usually, the nonlinear least-square problems involve more equations than the number of variables (#rows of $\boldsymbol{J}\gg$ #columns of $\boldsymbol{J}$), so the system is overdetermined. However, in the case of TopOpt, there are usually many more variables than equations (#rows of $\boldsymbol{J}\ll$ #columns of $\boldsymbol{J}$), making the system underdetermined. Therefore, the matrix $\boldsymbol{J}^{T}\boldsymbol{J}$ is large and rank-deficient, making the regular LMA formulation of Eq. (18) expensive and sometimes (when $\lambda_{i}\ll 1$) even unstable to solve.
@@ C.1 Dielectric 1D-layered stack [81]
[…]
@@
To calculate the gradient with respect to the layer thicknesses $\{d_{k}\}$, it is easy to calculate $dT/dd_{k}$, as only the element $T^{(k)}$ in Eq. (22) depends on $d_{k}$. Note that, in this case of transfer matrices, the adjoint method for calculating the derivative is essentially equivalent to saving all intermediate matrices $T^{(1...k-1)}$ during the forward k=1Kk=1...K matrix multiplications in Eq. (22), and then calculating $dT/dd_{k}$ backwards for k=K1k=K...1 and evaluating only once each matrix $T^{(k+1...K)}$ [9].
@@ C.2 Electronic $LC$-ladder circuit [34]
The $\mathcal{ABCD}$ matrix $\mathcal{T}$ of a two-port electronic circuit is defined as $\begin{pmatrix}v_{1}\\ i_{1}\end{pmatrix}=\begin{pmatrix}\mathcal{A}&\mathcal{B}\\ \mathcal{C}&\mathcal{D}\end{pmatrix}\begin{pmatrix}v_{2}\\ -i_{2}\end{pmatrix}$. Note that, although this $\mathcal{T}$ matrix relates voltage/current at port 1 to those at port 2, thus can also be cascaded, it is not the system transfer matrix $T$, which relates power-normalized inputs/outputs and requires knowledge of the input/output impedances. For an $LC$-ladder circuit, consisting of alternating series LLCC and shunt parallel $L\parallel C$ sections, the respective $\mathcal{ABCD}$ matrices are
[…]